OpenStudy (anonymous):

Solve each equation by factoring: 5r^2-44r+120=-30+11r 35k^2-22k+7=4

OpenStudy (tyteen4a03):

First question: 5r^2-44r+120=-30+11r can be transformed into: \[5r^2-55r+150=0\] Take out the 5 as a common factor before factoring, Second question: 35k^2-22k+7=4 can be transformed into: \[35k^2-22k+3\] If you see an equation that might be capable of being in the format \[ax^2 + bx + c = 0\], don't hesitate to try to turn the current equation into this format, then use either the Quadratic Equation or factoring to solve the equation.